# Reinforcement Learning Pair Trading: A Dynamic Scaling Approach

### v1

@@ 4.1 Pair Formation
Pairs are selected based on two criteria — correlation and cointegration. The widely adopted Pearson’s correlation [Perlin, 2007, Do & Faff, 2010] is given by
$\rho_{X,Y}=\frac{\mathrm{cov}(X,Y)}{\sigma_{X}\sigma_{Y}}$ (4)
where $\rho_{X,Y}$ is the correlation coefficient between assets $X$ and $Y$, $\mathrm{cov}(X,Y)$ is the covariance of $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively. The Engle-Granger cointegration test [Burgess, 2003, Dunis & Ho, 2005] involves two steps. First, the linear regression is performed:
$Y_{t}=\alpha+\beta X_{t}+\epsilon_{t}$ (5)
where $Y_{t}$ and $X_{t}$ are the asset price series, $\alpha$ and $\beta$ are the regression coefficients, and $\epsilon_{t}$ is the residual term. The second step tests the residuals $\epsilon_{t}$ for stationarity using an Augmented Dickey-Fuller (ADF)[Dickey & Fuller, 1979] test. The ADF test regression is given in Equation 6.
$\Delta\epsilon_{t}=\gamma\epsilon_{t-1}+\sum_{i=1}^{p}\delta_{i}\Delta\epsilon_{t-i}+\nu_{t}$ (6)
where $\Delta\epsilon_{t}$ is the first difference of the residuals, $\gamma$ is the coefficient to be tested for stationarity, $p$ is the number of lagged difference terms included, and $\nu_{t}$ is the error term. If $\gamma$ is significantly different from zero, the residuals are stationary, indicating cointegration.
A moving window is slid over historical pricing data (Figure 3). Averaged correlation and cointegration batches are used in the selection phase to ensure that the assets selected have a strong long-term statistical relationship.
Figure 3: Window-size cut for correlation and cointegration test
@@ 4.2 Spread Calculation
The second step is a moving window mechanism to capture the spread movement (Figure 2 Step-2). Spread $\epsilon_{t}$ is calculated at every selected trading interval (e.g., every five minutes) is the error term $s$ from a regression between the two prices $p_{i}$ and $p_{j}$, which follows normal distribution with mean of 0 and standard deviation of $\sigma$ (Equation 7):
$p_{i}=\beta_{0}+\beta_{1}\times p_{j}+s_{i}\sim N(0,\sigma^{2})$ (7)
We normalise the spread with $z$-score (Equation 8) to scale the spread into constant mean and standard deviation. The mean of the spread in the sliding window is represented as $\bar{s}$:
$Z=\frac{{s-\bar{s}}}{{\sigma_{s}}}$ (8)
@@ 4.3 Parameter Selection
Gatev et al. [2006] adopted 2 times standard deviation as the open threshold and the deviation crossing point as the close threshold (Figure 1). However, the thresholds ought to vary with market conditions. Hence, a window-sliding mechanism is incorporated to reflect the heterogeneity of the pricing variance [Mandelbrot, 1967].
Three parameters to be explored are 〈Window Size, Open Threshold, Close Threshold⟩. Window size is the number of historical samples in the moving window. Thresholds are highly linked to market conditions. Excessively wide thresholds suit more volatile markets, and conservatively narrow thresholds show result in smaller but steadier wins. The combination of parameters of the highest profitability 〈Window Size, Open Threshold, Close Threshold⟩ are selected from a search pool through grid search in practice.
@@ 4.4 Reinforcement Learning Pair Trading
After we run the grid search of window-sliding pair trading, the next problem concerns “when” and “how much” to trade. The pair trading results from following pre-set rules (à la Gatev et.al. 2006) are obtainable using window-sliding pair trading. However, we want to know if RL produces better investment decisions than blindly following the rules. Therefore, the most profitable parameter combination is passed onto further RL-based pair trading so that we can compare the results between RL-based pair trading and non-RL pair trading.
@@ 4.4.1 Observation Space
Observation space stands for the information an RL agent observes. The agent observes the market information to make decisions. The observations adopted for our RL environment is the following tuple: 〈Position, Spread, Zone⟩.
  • 1.
    Position $\in[-1,1]$: Position stands for the current portfolio value. Position is a percentage measuring the direction of investment (c.f. Figure 4). Assuming that we do not use leverage, holding a long leg with 70% portfolio value gives Position = 0.7. Holding a short leg with 30% portfolio value gives Position = -0.3. Position 0 means we only hold cash.
    Figure 4: The value of position observation based on investment
  • 2.
    Spread $\in\mathbb{R}$: It represents how much the current spread has deviated from the mean (Section 4.2).
  • 3.
    Zone $\in\{\text{Zones}\}$: Zone is an important indicator that comes from the comparison between the $z$-score with the thresholds for signals (Figure 1(b)). Traditional pair trading [Gatev et al., 2006, Yang & Malik, 2024] takes the zone as the direct trading signal (Figure 1(c)). However, in RL-based pair trading, the zone is an observation for the RL agent to make better decisions.
@@ 4.4.2 Action Space
Since the pair trading technique is a relatively low-risk strategy, most applications invest with a fixed amount or the complete portfolio value [Burgess, 2003, Perlin, 2009, Huck, 2010]. Meanwhile, it is natural for an experienced trader to invest different amounts based on the quality of opportunities. Opportunities with a higher probability of success are worth more investment. Therefore, we investigate granting the RL agent not only the decision of when to invest but also the freedom to choose the investment amount.
Take $A\in[-1,1]$ as the action. Similar to the observation space (c.f. Figure 4), action ranges from -1 to 1, representing the investment as a percentage of portfolio value to the long leg and short leg direction. Investing 50% of the portfolio as a long leg means A = 0.5. Investing 20% of the portfolio in short leg means A= -0.2.
In practice, we have to consider the relationship between the existing position and the next action. We classify the execution of action as 〈Open Position, Adjust Position, Close Position⟩.
  • 1.
@@
  • 3.
    Adjust Position happens when a previous position is open and the RL agent is wants to open another potion. For example, if the current position is 70% long-leg, and the new action is A=0.8, only the extra 10% shall be actioned.
@@ 4.4.3 Reward Shaping
The RL reward consists of three components: 〈Action Reward, Portfolio Reward, Transaction Punishment⟩.
  • 1.
    Portfolio Reward is the profit/loss from closing a position. The portfolio value $V_{p}$ updates only on position closing. If $V^{\prime}_{p}$ is the position value at the start of a trading period $p$. Then, upon closing the trade at the end of trading period $p$, the reward it calculated as shown in Equation (9).
    Profit $\text{Profit Reward}=V_{p}-V^{\prime}_{p}$ (9)
  • 2.
    Action Reward means the agent needs to be rewarded for taking a desired action in the corresponding zone. In general, the agent is free to decide on any action. However, we use Action Reward to encourage the agent to choose desired actions with Action Reward. It rewards the agent for making the desired action in certain zones (Table 1) with some freedom in neutral zones. The stronger the action reward, the more it resembles traditional pair trading.
    Zones Rewarding Behaviour
@@
    Table 1: Rewarding behaviours in zones
  • 3.
    Transaction Punishment is a negative reward for encouraging small adjustments instead of large changes in the position. The punishment is the difference between the action and position. If the current position in observation is $P$ and the action is $A$, the transaction punishment is (Equation 10):
    Transaction $\text{Transaction Punishment}=P-A$ (10)
@@ 5.1 Experimental Setup
We experiment with window-sliding pair trading and RL pair trading in the cryptocurrency market. The cryptocurrency market is famous for its volatility, easy access and 24/7 operating time.
@@ 5.1.1 Datasets
Figure 5: Prices of BTCEUR and BTCGBP
The application of our trading methodology is on Binance — the largest cryptocurrency market [^1] https://www.binance.com/en. For the best market liquidity, we picked Bitcoin - Fiat currencies under different trading intervals for pair trading. Pair formation criteria are based on Pearson’s correlation and augmented Engle-Granger two-step cointegration test (Section 4.1) for quote currencies that follow a similar trend against the base currency (Figure 5). The formation period is Oct-2023 to Nov-2023, and the test is in Dec-2023, with trading intervals of 1min (121,500 entries), 3min (40,500 entries), and 5min (24,300 entries), respectively. We exhaustively compared correlation and cointegration for the best pair (Table 2)  [^2] While calculating the Cointegration and Correlation, intervals with low volume trades are exempted from the calculation.. Though Binance has quite a few fiat currencies, only the US Dollar (USD), Great Britain Pounds (GBP), Euro (EUR) and Russian Ruble (RUB) display relatively strong liquidity. The pair with the strongest correlation and cointegration is BTCEUR and BTCGBP under 1min trading interval (2).
@@
Table 2: Correlation and cointegration of pair formation
Transaction cost in the experiment is set to 0.02% commission based on Binance’s fee scheme  [^3] https://www.binance.com/en/fee/futureFee. The transaction cost of 0.02% is a flat percentage charge for transactions in both directions. A pair trading leg, including long the first asset and short the second asset, is charged for both long and short actions.
@@ 5.1.2 Grid Search and Reinforcement Learning
Grid search is used to find a set of profitable parameters, including open/close threshold and window size during the training period (Oct-2023 to Nov-2023). Every iteration of the window-sliding pair trading experiments with one set of parameters (window size, open/close threshold) until exhaustion. The most profitable parameter set will be used to test traditional pair trading during the test period (Dec-2023) and also for testing the proposed RL-strategy.
The profitability in grid search is measured by Total Compound Return (RTOT), where $V_{p}$ and $V_{p}^{\prime}$ are the value of the portfolio at the beginning of the period and the end of the period and $t$ as the total length of the trading period (Equation 11):
$rtot=(V_{p}^{\prime}/V_{p})^{1/t}-1\times 100\%$ (11)
During the training period, the most profitable parameter set is〈open threshold = 1.8 $z$-score, close threshold = 0.4 $z$-score, window size = 900 intervals⟩. Some example results of the grid search are presented in Table 3.
@@
Table 3: Trading Parameters Tuning
The setup of RL-based pair trading relies on these parameters. The window size decides the retrospective length of the spread, and the thresholds decide the zones. Algorithms such as PPO and A2C are applicable to both discrete and continuous action spaces. Some algorithms, e.g. DQN, can only be used on discrete space, and DDPG is only applicable in a continuous space. Therefore, we adopt PPO, DQN and A2C on RL pair trading that decides the timing, and PPO, A2C,SAC on RL pair trading that decides both timing and investment quantity. The algorithms are adopted from the Baseline3 collection [Raffin et al., 2021].
@@ 5.1.3 Evaluation Metrics
Our main concern is the highest profitability in trading techniques. We care about the cumulative return, which is the total profit for the testing period, as well as the annualised return Compound Annual Growth Rate (CAGR). With $V(t_{0})$ as the initial state, $V(t_{n})$ as the final state, $t_{n}-t_{0}$ is period of trading in years, the CAGR is (Equation 12):
$\operatorname{CAGR}\left(t_{0},t_{n}\right)=\left(\frac{V\left(t_{n}\right)}{V\left(t_{0}\right)}\right)^{\frac{1}{t_{n}-t_{0}}}-1$ (12)
There are some popular indicators for distinguishing whether a strategy is profit-risk effective, eg. Sharpe Ratio. Sharpe Ratio [Sharpe, 1964] where $R_{p}$ is the return of the trading strategy, $R_{f}$ is the interest rate [^4] We adopt the Federal Reserve interest rate 5.5% as on 13 June 2024, and $\sigma_{p}$ is the standard deviation of the portfolio’s excess return (Equation 13):
Sharpe $\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (13)
We also care about the strategies’ activities, such as order count and win/loss ratio.
The indicators used for comparison are presented in Table 4.
@@
Avg Action P&L Average profit/loss per trade in Bitcoin

### v2

@@ 4.1 Pair Formation
Pairs are selected based on two criteria: correlation and cointegration. The widely adopted Pearson’s correlation (Perlin, 2007; Do and Faff, 2010) is given by
$\rho_{X,Y}=\frac{\mathrm{cov}(X,Y)}{\sigma_{X}\sigma_{Y}},$ (4)
where $\rho_{X,Y}$ is the correlation coefficient between assets $X$ and $Y$, $\mathrm{cov}(X,Y)$ is the covariance of $X$ and $Y$, and $\sigma_{X}$ and $\sigma_{Y}$ are the standard deviations of $X$ and $Y$, respectively. The Engle–Granger cointegration test (Burgess, 2003; Dunis and Ho, 2005) involves two steps. First, the linear regression is performed:
$Y_{t}=\alpha+\beta X_{t}+\epsilon_{t},$ (5)
where $Y_{t}$ and $X_{t}$ are the asset price series, $\alpha$ and $\beta$ are the regression coefficients, and $\epsilon_{t}$ is the residual term. The second step tests the residuals $\epsilon_{t}$ for stationarity using an Augmented Dickey-Fuller (ADF)(Dickey and Fuller, 1979) test. The ADF test regression is given in Equation (6):
$\Delta\epsilon_{t}=\gamma\epsilon_{t-1}+\sum_{i=1}^{p}\delta_{i}\Delta\epsilon_{t-i}+\nu_{t},$ (6)
where $\Delta\epsilon_{t}$ is the first difference of the residuals, $\gamma$ is the coefficient to be tested for stationarity, $p$ is the number of lagged difference terms included, and $\nu_{t}$ is the error term. If $\gamma$ is significantly different from zero, the residuals are stationary, indicating co-integration.
A moving window is applied to historical pricing data, as shown in Figure 3. In this figure, the blue line represents the historical prices, while the dashed boxes illustrate the moving window. During the selection phase, averaged correlation and co-integration batches are employed to ensure that the selected assets exhibit a strong, long-term statistical relationship.
Figure 3: Window-size cut for correlation and co-integration testing.
@@ 4.2 Spread Calculation
The second step is a moving window mechanism to capture the spread movement (Figure 2 Step-2). Spread $\epsilon_{t}$ is calculated at every selected trading interval (e.g., every five minutes) and is the error term $s$ from a regression between the two prices $p_{i}$ and $p_{j}$. $\beta_{0}$ and $\beta_{1}$ are the intercept and slope, respectively, which follow a normal distribution with a mean of 0 and a standard deviation of $\sigma$ Equation (7):
$p_{i}=\beta_{0}+\beta_{1}\times p_{j}+s_{i}\sim N(0,\sigma^{2}).$ (7)
We normalize the spread with $z$-score Equation (8) to scale the spread into constant mean and standard deviation. The mean of the spread in the sliding window is represented as $\bar{s}$:
$Z=\frac{{s-\bar{s}}}{{\sigma_{s}}}.$ (8)
@@ 4.3 Parameter Selection
Three parameters to be explored are window size, open threshold, and close threshold. Window size $\in\mathbb{Z}^{+}$ is the number of historical samples in the moving window. Thresholds $\in\mathbb{Q}^{+}$ are the entry and exit signals of trading actions that are highly linked to market conditions.
Excessively wide thresholds suit more volatile markets, and conservatively narrow thresholds result in smaller but steadier wins. The combination of parameters of the highest profitability 〈Window Size , Open Threshold , Close Threshold ⟩ are selected from a search pool through a grid search in practice. Gatev et al. (2006) adopted 2 times the standard deviation as the open threshold and the deviation crossing point as the close threshold (Figure 1). However, on the one hand, the parameters should vary with the arbitrage instruments as well as the market condition. Hence, the window-sliding mechanism is incorporated to reflect the heterogeneity of the pricing variance (Mandelbrot, 1967).
@@ 4.4 Reinforcement Learning Pair Trading
After we run the grid search of window-sliding pair trading, the next problem concerns “when” and “how much” to trade. Pair trading results from following pre-set rules (à la Gatev et al. 2006), which are obtainable using window-sliding pair trading. However, we want to know if RL produces better investment decisions than blindly following the rules. Therefore, the most profitable parameter combination is passed onto further RL-based pair trading so that we can compare the results between RL-based pair trading and non-RL pair trading.
@@ 4.4.1 Observation Space
Observation space stands for the information an RL agent observes. The agent observes market information to make decisions. The observations adopted for our RL environment are as follows: 〈Position, Spread, Zone⟩.
  • Position $\in[-1,1]$: Position stands for the current portfolio value. Position is a percentage measuring the direction of investment (c.f. Figure 4). As in pair trading, we define longing the first asset with shorting the second asset as ‘holding a long leg’ and the other way around as ‘holding a short leg’. Assuming that we do not use leverage, holding a long leg with a 70% portfolio value gives Position = 0.7. Holding a short leg with a 30% portfolio value gives Position = --0.3. Position 0 means we only hold cash.
  • Spread $\in\mathbb{R}$: This represents how much the current spread has deviated from the mean (Section 4.2).
  • Zone $\in\{\text{Zones}\}$: Zone is an important indicator that comes from the comparison between the $z$-score with the thresholds for signals (Figure 1b). Traditional pair trading (Gatev et al., 2006; Yang and Malik, 2024) takes the zone as the direct trading signal. However, in RL-based pair trading, the zone is an observation for the RL agent to make better decisions.
Figure 4: The value of position observation based on investment.
@@ 4.4.2 Action Space
Since the pair trading technique is a relatively low-risk strategy, most applications invest with a fixed amount or the complete portfolio value (Burgess, 2003; Perlin, 2009; Huck, 2010). Meanwhile, it is natural for an experienced trader to invest different amounts based on the quality of opportunities. Opportunities with a higher probability of success are worth more investment. Therefore, we investigate granting the RL agent not only the decision of when to invest but also the freedom to choose the investment amount.
Take $A\in[-1,1]$ as the action. Similar to the observation space (c.f. Figure 4), the action ranges from --1 to 1, representing the investment as a percentage of the portfolio value in the long leg and short leg directions. Investing 50% of the portfolio as a long leg means A = 0.5. Investing 20% of the portfolio in the short leg means A= --0.2.
In practice, we have to consider the relationship between the existing position and the next action. We classify the execution of action as open position , adjust position , or close position .
@@
  • Adjust position happens when a previous position is open, and the RL agent wants to open another position. For example, if the current position is a 70% long-leg and the new action is A = 0.8, only the extra 10% shall be actioned.
@@ 4.4.3 Reward Shaping
The RL reward consists of three components: action reward , portfolio reward , and transaction punishment .
  • Portfolio reward is the profit/loss from closing a position. The portfolio value $V_{p}$ only updates when the position closes. $V^{\prime}_{p}$ is the position value at the start of a trading period $p$. Then, upon closing the trade at the end of trading period $p$, the reward it calculated, as shown in Equation (9).
    Profit $\text{Profit Reward}=V_{p}-V^{\prime}_{p}.$ (9)
  • Action reward means the agent needs to be rewarded for taking a desired action in the corresponding zone. In general, the agent is free to decide on any action. However, we use action reward to encourage the agent to choose desired actions. It rewards the agent for making a desired action in certain zones (Table 1) with some freedom in neutral zones. The stronger the action reward, the more it resembles traditional pair trading.
  • Transaction punishment is a negative reward for encouraging small adjustments instead of large changes in the position. The punishment is the difference between the action and position. If the current position in observation is $P$ and the action is $A$, the transaction punishment is Equation (10):
    Transaction $\text{Transaction Punishment}=P-A.$ (10)
Table 1: Rewarding behaviors in zones.
                  Zones                   Rewarding Behavior
@@
4.5 Dynamic Agents
Trading agents in RL environments operate by taking actions within a defined action space based on observations from the state space. This section details the design of two dynamic agent settings, RL1 and RL2, each tailored to address specific aspects of pair trading. These agents leverage the Markov Decision Process (MDP) framework to model the complexities of financial markets, enabling adaptive decision-making in volatile conditions.
In the first setting, RL1 is tasked with optimizing trade timing and directionality. Pair trading is modeled as a Markov Decision Process (MDP), represented as $(\mathcal{S}_{1},\mathcal{A}_{1},\mathcal{T}_{1},r_{1},\gamma_{1})$:
  • $\mathcal{S}_{1}$ represents the state space, including normalized price spreads, historical z-scores, and zone indicator.
  • $\mathcal{A}_{1}$ defines a discrete action space consisting of three possible actions: initiating a long–short position, closing existing positions, or initiating a short–long position. This allows the agent to determine the optimal direction and timing for trades.
  • $r_{1}(s,a)$ is the reward function, defined as follows:
    $r_{1}(s_{t},a_{t})=\begin{cases}\Delta P_{t}-c,&\text{if a trade is executed;}\\ 0,&\text{if no trade is executed,}\end{cases}$
    where $\Delta P_{t}$ represents the profit or loss from the trade, and $c$ is the transaction cost. The reward function penalizes the agent for transaction costs while directly linking rewards to trade profitability.
  • The goal is to maximize the cumulative discounted reward:
    $R_{1}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{1}^{t}r_{1}(s_{t},a_{t})\right].$
RL2 extends RL1 by shifting focus from trade timing to determining the investment quantity for a given trade opportunity. It models pair trading as an MDP defined by $(\mathcal{S}_{2},\mathcal{A}_{2},\mathcal{T}_{2},r_{2},\gamma_{2})$:
  • $\mathcal{S}_{2}$ is the state space, which is identical to RL1.
  • $\mathcal{A}_{2}=[-1,1]$, where the continuous value represents the investment quantity. Here, 00 stands for no involvement, positive values represent buying, and negative values represent selling.
  • $r_{2}(s,a)$ is the reward function, defined as follows:
    $r_{2}(s_{t},a_{t})=\Delta P_{t}\cdot a_{t}-c(|a_{t}|),$
    where $c(|a_{t}|)$ represents transaction costs proportional to the absolute investment size $|a_{t}|$. This reward structure incentivizes the agent to optimize both the direction and magnitude of its investment.
  • The objective is to maximize the cumulative discounted reward:
    $R_{2}=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{2}^{t}r_{2}(s_{t},a_{t})\right].$
The primary differences between RL1 and RL2 lie in their action spaces and reward functions. RL1 operates with a discrete action space and focuses on optimizing directional timing and trade management. In contrast, RL2 uses a continuous action space $\mathcal{A}_{2}=[-1,1]$, enabling it to adjust investment sizes dynamically. The environments are designed to guide the agents by rewarding profitable actions and penalizing costly ones, encouraging effective decision-making for timing and quantity. The exact mechanisms driving these decisions are embedded within the neural network, shaped by the agent’s interactions with the environment.
@@ 5.1 Experimental Setup
We experiment with window-sliding pair trading and RL pair trading in the cryptocurrency market. The cryptocurrency market is famous for its volatility, easy access, and 24/7 operating time. Data preprocessing follows two steps. Firstly, the evaluation of the correlations among trading pairs for selection of arbitrage candidates; next, searching through the combination of thresholds and retrospective periods for suitable signals.
@@ 5.1.1 Datasets
[…]
Figure 5: Prices of BTCEUR and BTCGBP.
Table 2: Correlation and co-integration of pair formation.
@@
[…]
@@ 5.1.2 Grid Search and Reinforcement Learning
A grid search is used to find a set of profitable parameters, including the open/close threshold and window size during the training period (October 2023 to November 2023). Every iteration of the window-sliding pair trading experiments has one set of parameters (window size, open/close threshold) until exhaustion. We start off the exploration based on the experiential estimation of the asset characteristics. The most profitable parameter set will be used to test traditional pair trading during the test period (December 2023) and also for testing the proposed RL strategy.
The profitability in conducting a grid search is measured by the Total Compound Return (RTOT), where $V_{p}$ and $V_{p}^{\prime}$ are the values of the portfolio at the beginning of the period and the end of the period, and $t$ is the total length of the trading period in Equation (11):
$rtot=(V_{p}^{\prime}/V_{p})^{1/t}-1\times 100\%.$ (11)
During the training period, the most profitable parameter set is open threshold = 1.8 $z$-score, close threshold = 0.4 $z$-score, and window size = 900 intervals. Some example results of the grid search are presented in Table 3.
Table 3: Trading parameter tuning.
@@
The setup of RL-based pair trading relies on these parameters. The window size decides the retrospective length of the spread, and the thresholds decide the zones. Algorithms such as PPO and A2C are applicable to both discrete and continuous action spaces. Some algorithms, e.g., DQN, can only be used in discrete space, and DDPG is only applicable in a continuous space. Therefore, we adopt PPO, DQN, and A2C in RL pair trading, which decide the timing, and PPO, A2C, and SAC in RL pair trading, which decide both the timing and investment quantity. The algorithms are adopted from the Baseline3 collection (Raffin et al., 2021).
@@ 5.1.3 Evaluation Metrics
Our main concern is the highest profitability in trading techniques. We care about the cumulative return, which is the total profit for the testing period, as well as the annualized return Compound Annual Growth Rate (CAGR). With $V(t_{0})$ as the initial state, $V(t_{n})$ as the final state, and $t_{n}-t_{0}$ as the period of trading in years, the CAGR is Equation (12):
$\operatorname{CAGR}\left(t_{0},t_{n}\right)=\left(\frac{V\left(t_{n}\right)}{V\left(t_{0}\right)}\right)^{\frac{1}{t_{n}-t_{0}}}-1.$ (12)
[…]
Sharpe $\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}.$ (13)
We also care about the strategies’ activities, such as the order count and win/loss ratio. The indicators used for comparison are presented in Table 4.
Table 4: Descriptive table of evaluation metrics.
@@
          Avg Action Profit/Loss           Average profit/loss per trade in Bitcoin